A factory floor has 777 square meters. If its width is 24 meters, find the length. I need help with this problem
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

A factory floor has 777 square meters. If its width is 24 meters, find the length.
-------
Area = L*W
777 = L*24

Answer by fractalier(6550)

Given that the floor is rectangular, you can find the length from
Area A = LW which gives us
L = A/W = 777/24 = 32.375 meters